On January 1, 2021, Splash City issues $430,000 of 8% bonds, due in 15 years, with interest payable semiannually on June 30 and December 31 each year.

Assuming the market interest rate on the issue date is 9%, the bonds will issue at $394,979.

I need help completing the first three rows of the amortization table.
Answer #1

Amortization schedule

Date Interest expense Interest paid Amortization Carrying value
Jan 1 394979
June 30 394979*4.5% = 17774 17200 574 395553
Dec 31 395553*4.5% = 17800 17200 600 396153
June 30 396153*4.5% = 17827 17200 627 396780